OpenStudy (anonymous):

solve the equation for 2^x-5=3

OpenStudy (anonymous):

2^x-5=3 \[2^x=5+3\] \[2^x = 8\] x=3 do you got it @JosephineM

OpenStudy (anonymous):

yes thank you so much! I totally understand now!

OpenStudy (anonymous):

:) :)